# Parametric Expensive Multi-Objective Optimization via Generative Solution Modeling

### v3

@@ 4 Results
We provide empirical experiments demonstrating our method’s ability to build effective inverse models for P-EMOP. Our experiments verify three key claims: (1) PMT-MOBO achieves faster convergence by exploiting inter-task synergies, (2) conditional generative models further improve solution quality, and (3) the built inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ can directly predict optimized solutions for unseen EMOP tasks without additional evaluations. We evaluate all methods using the hypervolume indicator [67] across both seen and unseen EMOP tasks.
@@
We evaluate our methods on classical synthetic benchmarks: DTLZ-1, DTLZ-2, and DTLZ-3 [13], and real-world problems including lamp generation, solar rooftop generation [35], magnetic sifter design [33], and UAV controller design [52]. We modify these base EMOP problems so that each benchmark defines a family of related EMOPs, where $\boldsymbol{\theta}$ represents operational conditions (e.g., focal heights of lamps, building orientations of rooftops, and physical configurations of UAVs). More problem details are provided in the Appendix B.
@@
Our empirical studies include state-of-the-art single-task MOBO methods: ST-MOBO [41], qPAREGO [1], qEHVI [11], PSL-MOBO [32], and CDM-PSL [30], a diffusion model based MOBO variant. Moreover, to evaluate performance in utilizing inter-task synergies, we compare against multi-task MOBO methods: OCMOBO [6] and PPSL-MOBO [8], a parametric extension of PSL-MOBO [32]. Finally, we evaluate variants of our approaches that serve as ablation studies: PMT-MOBO (acquisition-driven only), PMT-MOBO-VAE, and PMT-MOBO-DDPM (full framework with different instantiated generative models).
@@
Figure 2, Table 1 and Table 2 demonstrate our framework’s effectiveness on synthetic and real-world benchmarks. PMT-MOBO-VAE and PMT-MOBO-DDPM generally show superior or comparable results compared to state-of-the-art single-task and multi-task MOBO baselines in discovering high-quality Pareto solutions across distinct tasks. The exception in magnetic sifter motivates us to further explore the integration of our methods with EHVI based acquisition functions so that the proposed framework can be more flexibly adapt to diverse applications. Table 2 further shows that our inverse models generalize well to unseen EMOP tasks, outperforming the single-task and multi-task inverse models constructed by PSL-MOBO and PPSL-MOBO. To evaluate this generalization capability, we extract the trained inverse models after optimization and apply them to $W=100$ completely new tasks. For each unseen task, we query the inverse model with $S$ uniformly sampled preference vectors ($S=100$ for bi-objective, $S=1000$ for tri-objective) to generate predicted solutions, then compute the hypervolume of the resulting solution set.
@@
Table 1 summarizes the hypervolume results across all benchmarks. Wilcoxon signed-rank tests (at 90% confidence) demonstrate that our proposed methods significantly outperform the baselines in most scenarios. Specifically, against the seven competing methods across all tasks (49 pairwise comparisons), PMT-MOBO-VAE attains a win/tie/loss record of 38/5/6, effectively dominating difficult tasks such as LAMP and SOLAR. The PMT-MOBO-DDPM further enhances optimization capability, achieving a record of 48/0/1, statistically surpassing all baselines in every benchmark except for the low-dimensional magnetic sifter design.
@@
Method DTLZ-1 DTLZ-2 DTLZ-3 LAMP SOLAR UAV
PSL-MOBO 2.82e+4 (1.78e+3) 1.90e+0 (1.30e-1) 2.39e+4 (3.22e+3)+- 5.01e-1 (7.32e-2) 3.54e-1 (4.06e-2) 4.12e-1 (7.30e-3)
PPSL-MOBO 2.24e+4 (1.10e+2) 2.98e+0 (3.56e-2) 2.67e+4 (6.08e+2)+- 6.12e-1 (1.05e-1)≈- 4.86e-1 (1.11e-1) 5.24e-1 (1.71e-1)
PMT-MOBO-VAE 2.88e+4 (2.57e+3) 3.10e+0 (2.99e-2) 1.78e+4 (7.28e+3) 6.31e-1 (7.33e-2) 7.93e-1 (1.56e-2) 6.08e-1 (1.88e-2)
PMT-MOBO-DDPM 3.47e+4 (4.85e+2) 3.13e+0 (1.41e-2) 3.20e+4 (1.98e+3) 6.70e-1 (5.82e-2) 7.65e-1 (1.30e-2) 6.05e-1 (2.44e-2)
Table 2: Inverse Model Generalization Performance on Unseen EMOP Tasks. Statistical significance is assessed using the Wilcoxon signed-rank test at the 90% confidence level. Superscripts $^{s_{1}s_{2}}$ indicate that the competitor is significantly worse (-), significantly better (++), or statistically similar ($\approx$) relative to PMT-MOBO-VAE ($s_{1}$) and PMT-MOBO-DDPM ($s_{2}$).
@@
To validate the necessity of our alternating optimization strategy, we compared the full framework against decoupled baselines (pure acquisition search and pure generative sampling). The results, detailed in Appendix D, confirm that neither component is sufficient in isolation. Specifically, generative sampling enhances the diversity of candidate generation, while the alternating acquisition search lowers the variance inherent in pure generative sampling, resulting in robust convergence. Moreover, figure 2 demonstrates the superior performance of PMT-MOBO-VAE and PMT-MOBO-DDPM compared to PMT-MOBO, also highlighting the value of integrating generative modeling.
Furthermore, our sensitivity analysis on the elite ratio parameter $Q$ reveals a critical trade-off: too large a $Q$ reduces selective pressure, pushing the generative model to learn low-quality solutions, whereas too small a $Q$ incurs large variance and provides insufficient data to train a well-behaved model. Empirically, $Q=10\%$ achieves the optimal balance across both VAE and DDPM variants (see Tables 6 and  7 in Appendix.D).
@@ Appendix B Benchmark Definitions
In Table 3, we specify the dimension of solution and task spaces for each benchmark problem as well as the number of objectives. Note that, since the problem formulation we focus on is expensive multi-objective optimization problems, too large a solution or task space might easily incur a dimensionality curse for those surrogate models and generative models. Therefore, in this paper, we mainly focus on solving low-to-moderate P-EMOPs. And we leave solving high-dimensional P-EMOPs as a future challenging direction here.
@@ B.2 Real-World Benchmarks
Appendix C Experimental Setup
The source code link will be made public upon acceptance.
C.1 Generative Model Training
For the DDPM, we use a simple MLP architecture that takes the concatenated input $\mathbf{x},t,\boldsymbol{c}]$, where $t$ is the timestep, and $\mathbf{c}$ is the conditioning vector. The network consists of 4 fully connected layers with hidden dimension 128, following the structure of [28]. We employ a linear noise schedule with $\beta$ values ranging from 0.0001 to 0.02 over 1000 time steps. The model is trained to predict the noise $\boldsymbol{\epsilon}$ added during the forward diffusion process using standard MSE loss. During training, we randomly sample timesteps $t\sim\text{Uniform}(0,T-1)$ and compute the loss $\mathcal{L}=\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{c})\|_{2}^{2}$. For sampling, we use the standard DDPM reverse process, starting from Gaussian noise and iteratively denoising over all timesteps. We apply gradient clipping with a maximum norm of 1.0 for training stability and clamp the generated samples to the $[0,1]$ range to handle boundary constraints.
C.2 GP Model Training
where $\kappa_{\text{dec}}$ is an isotropic RBF kernel for decision variables with lengthscale constrained to the interval $[0.1,2.5]$, and $\kappa_{\text{task}}$ is an ARD RBF kernel for task parameters, allowing each task dimension to have its own lengthscale parameter. The overall covariance function is wrapped by a scale kernel to scale the product of the two kernels. This composite kernel is partially inspired by the kernel function in [6]. The GP-based source codes are partially supported by GPytorch [16].
Appendix D Detailed Experimental Results
D.1 Ablation Studies
Variants DTLZ-1 DTLZ-2 DTLZ-3
Variants DTLZ-1 DTLZ-2 DTLZ-3
Tables 4 and 5 present the ablation results for VAE and DDPM integrations, respectively. We isolate the impact of the alternating loop by comparing the full method against two decoupled baselines: pure acquisition search (PMT-MOBO w/o GENERATIVE) and pure generative sampling (PLAIN-VAE/DDPM). The results for the VAE variant (Table 4) clearly illustrate the limitations of isolated components. PLAIN-VAE exhibits a high standard deviation across benchmarks (e.g., $7.56\text{e}3$ on DTLZ-1), indicating that while pure generative sampling can occasionally "get lucky" and hit high-quality solutions, it suffers from inherent stochasticity during the optimization process and lacks the precision to consistently refine the Pareto front. Conversely, pure acquisition search often converges prematurely to local optima, failing to explore promising but unseen regions of the design space. By coupling these components, PMT-MOBO-VAE leverages generative sampling to enhance candidate diversity while utilizing the alternating acquisition search to lower the variance of pure sampling. This synergy results in statistically superior convergence and stability across all tested problems.
@@ D.1 Ablation Studies
D.2 Sensitivity Analysis

### v4

@@ 4 Results
We provide empirical experiments demonstrating our method’s ability to build effective inverse models for P-EMOPs. Our experiments verify three key claims: (1) PMT-MOBO achieves faster convergence by exploiting inter-task synergies, (2) conditional generative models further improve solution quality, and (3) the built inverse model $\mathcal{M}(\boldsymbol{\theta},\boldsymbol{\lambda})$ can directly predict optimized solutions for unseen EMOP tasks without additional evaluations. We evaluate all methods using the hypervolume indicator [61] across both seen and unseen EMOP tasks.
@@
We evaluate our methods on classical synthetic benchmarks: DTLZ-1, DTLZ-2, and DTLZ-3 [12], and real-world problems including lamp generation, solar rooftop generation [33], magnetic sifter design [31], and UAV controller design [48]. We modify these base EMOP problems so that each benchmark defines a family of related EMOPs, where $\boldsymbol{\theta}$ represents operational conditions (e.g., physical configurations of UAVs). More problem details are provided in the Appendix B.
@@
Our empirical studies include state-of-the-art single-task MOBO methods: ST-MOBO [39], qPAREGO [1], qEHVI [11], PSL-MOBO [30], and CDM-PSL [29], a diffusion model based MOBO variant. Moreover, to evaluate performance in utilizing inter-task synergies, we compare against multi-task MOBO methods: OCMOBO [6] and PPSL-MOBO [8]. Finally, we evaluate variants of our approaches that serve as ablation studies: PMT-MOBO (acquisition-driven only), PMT-MOBO-VAE, and PMT-MOBO-DDPM (full framework with different generative models).
@@
Figure 2, Table 5 and Table 1 demonstrate our framework’s effectiveness on synthetic and real-world benchmarks. PMT-MOBO-VAE and PMT-MOBO-DDPM generally show superior or comparable results compared to state-of-the-art single-task and multi-task MOBO baselines in discovering high-quality Pareto solutions across distinct tasks. The exception in magnetic sifter motivates us to further explore the integration of our methods with EHVI based acquisition functions so that the proposed framework can be more flexibly adapt to diverse applications. Table 1 further shows that our inverse models generalize well to unseen EMOP tasks, outperforming the single-task and multi-task inverse models constructed by PSL-MOBO and PPSL-MOBO. With task-aware kernel and the alternating optimization loop, our PMT-MOBO-DDPM’s inverse model can also outperform the diffusion model of CDM-PSL across all the problems. To evaluate this generalization capability, we extract the trained inverse models after optimization and apply them to $W=100$ completely new tasks. For each unseen task, we query the inverse model with $S$ uniformly sampled preference vectors ($S=100$ for bi-objective, $S=1000$ for tri-objective) to generate predicted solutions, then compute the hypervolume of the resulting solution set.
@@
Figure 2 and Table 5 in Appendix D.1 summarize the hypervolume results across all benchmarks. Wilcoxon signed-rank tests (at 95% confidence) demonstrate that our proposed methods significantly outperform the baselines in most scenarios. Specifically, against the eight competing methods across all tasks (56 pairwise comparisons), PMT-MOBO-VAE attains a win/tie/loss record of 45/8/3, effectively dominating difficult tasks such as LAMP and SOLAR. The PMT-MOBO-DDPM further enhances optimization capability, achieving a record of 55/0/1, statistically surpassing all baselines in every benchmark except for the low-dimensional magnetic sifter design.
@@
To validate the necessity of our alternating optimization strategy, we compared the full framework against decoupled baselines (pure acquisition search and pure generative sampling). The results, detailed in Appendix D.2, confirm that neither component is sufficient in isolation. Specifically, generative sampling enhances the diversity of candidate generation, while the alternating acquisition search lowers the variance inherent in pure generative sampling, resulting in robust convergence. Figure 2 also demonstrates the superior performance of PMT-MOBO-VAE and PMT-MOBO-DDPM compared to PMT-MOBO, highlighting the value of integrating generative modeling.
Furthermore, our sensitivity analysis on the elite ratio parameter $Q$ reveals a critical trade-off: too large a $Q$ reduces selective pressure, pushing the generative model to learn low-quality solutions, whereas too small a $Q$ incurs large variance and provides insufficient data to train a well-behaved model. Empirically, $Q=10\%$ achieves the optimal balance across both VAE and DDPM variants (see Tables 6 and 7 in Appendix D.3).
@@ Appendix A Proofs of Theorem 2
Remark
The theoretical guarantee established in Theorem 2 relies on the covariance reduction property of the task-aware kernel structure, where task kernel reduces to single-task kernel when the solution pair belongs to the same tasks. We note that our composite kernel, defined as $\tilde{\kappa}((x,\theta),(x^{\prime},\theta^{\prime}))=\kappa_{dec}(x,x^{\prime})\cdot\kappa_{task}(\theta,\theta^{\prime})$, aligns with this theoretical assumption and represents a widely applied formulation in multi-task Gaussian Process surrogates [6]. The validity of this tighter regret bound is empirically substantiated by the superior convergence rates observed across our benchmarks, as detailed in Table 5 and Figure 2. Crucially, the theoretical derivation of Theorem 2 introduces no additional requirements beyond those already established in Assumption 1. Since the noisy observations remain sub-Gaussian and the scalarization function remains monotone and Lipschitz, the structural validity of the original regret upper bound in Theorem 1 is preserved. Theorem 2 thereby demonstrates that under these identical regularity conditions, the multi-task covariance reduction property strictly decreases the Maximum Information Gain (MIG), the core governing term of the regret bound, thereby tightening the bound without necessitating more restrictive assumptions.
Furthermore, we explicitly acknowledge that while the current theoretical framework guarantees faster convergence for the acquisition-driven search via Task-Aware GPs, it does not yet derive a formal regret bound for the alternating generative solution sampling. The generative step acts as a heuristic proposal distribution to help mitigate the bias of pure acquisition search and escape local optima, bolstered by our empirical studies and ablation studies. Extending the theoretical analysis to unify both the multi-objective Bayesian optimization bounds and the generative sampling dynamics remains a non-trivial challenge. It is left as a potential direction for future work.
@@ Appendix B Benchmark Definitions
Variants DTLZ-1 DTLZ-2 DTLZ-3
Variants DTLZ-1 DTLZ-2 DTLZ-3
In Table 2, we specify the dimension of solution and task spaces for each benchmark problem as well as the number of objectives. Note that, since the problem formulation we focus on is expensive multi-objective optimization problems, too large a solution or task space might easily incur a dimensionality curse for those surrogate models and generative models. Therefore, in this paper, we mainly focus on solving low-to-moderate P-EMOPs. We leave solving high-dimensional P-EMOPs as a future challenging direction.
@@ B.2 Real-World Benchmarks
Method DTLZ-1 DTLZ-2 DTLZ-3 LAMP SOLAR MAGNETIC UAV
qEHVI 3.23e+04 (5.06e+01)≈- 2.33e+00 (6.20e-02) 2.66e+04 (1.85e-02)+- 6.14e-01 (3.00e-03) 6.07e-01 (1.29e-02) 2.18e-01 (1.15e-02)++ 4.89e-01 (2.60e-03)
CDM-PSL 1.59e+04 (4.15e+03) 2.43e+00 (5.04e-02) 7.76e+02 (9.41e+02) 5.85e-01 (1.39e-02) 6.49e-01 (1.97e-02) 1.00e-01 (1.82e-02) 3.89e-01 (5.70e-03)
PPSL-MOBO 3.20e+04 (1.72e+03)≈- 2.86e+00 (3.00e-01)≈- 2.71e+04 (7.89e+02)+- 6.09e-01 (1.61e-02) 6.35e-01 (1.15e-02) 9.41e-02 (1.69e-02) 5.85e-01 (3.19e-02)≈-
PMT-MOBO 1.71e+04 (4.70e+03) 2.77e+00 (5.50e-02) 3.78e+03 (5.40e+03) 5.83e-01 (1.22e-02) 7.10e-01 (1.90e-02) 1.73e-01 (2.99e-02)≈- 4.78e-01 (1.27e-02)
Table 5: Average Hypervolume results (mean $\pm$ std) across synthetic and real-world benchmarks. Statistical significance is assessed using the Wilcoxon signed-rank test at the 95% confidence level. Superscripts $^{s_{1}s_{2}}$ indicate that the competitor is significantly worse (-), significantly better (++), or statistically similar ($\approx$) relative to PMT-MOBO-VAE ($s_{1}$) and PMT-MOBO-DDPM ($s_{2}$).
@@
Appendix C Experimental Setup
The source code link will be made public upon acceptance.
C.1 Generative Model Training
For the DDPM, we use a simple MLP architecture that takes the concatenated input $\mathbf{x},t,\boldsymbol{c}]$, where $t$ is the timestep, and $\mathbf{c}$ is the conditioning vector. The network consists of 4 fully connected layers with hidden dimension 128, following the structure of [27]. We employ a linear noise schedule with $\beta$ values ranging from 0.0001 to 0.02 over 1000 time steps. The model is trained to predict the noise $\boldsymbol{\epsilon}$ added during the forward diffusion process using standard MSE loss. During training, we randomly sample timesteps $t\sim\text{Uniform}(0,T-1)$ and compute the loss $\mathcal{L}=\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\phi}(\mathbf{x}_{t},t,\mathbf{c})\|_{2}^{2}$. For sampling, we use the standard DDPM reverse process, starting from Gaussian noise and iteratively denoising over all timesteps. We apply gradient clipping with a maximum norm of 1.0 for training stability and clamp the generated samples to the $[0,1]$ range to handle boundary constraints. To accelerate the DDPM sampling process, we adopte the widely applied dpm-solver [32].
C.2 GP Model Training
where $\kappa_{\text{dec}}$ is an isotropic RBF kernel for decision variables with lengthscale constrained to the interval $[0.1,2.5]$, and $\kappa_{\text{task}}$ is an ARD RBF kernel for task parameters, allowing each task dimension to have its own lengthscale parameter. The overall covariance function is wrapped by a scale kernel to scale the product of the two kernels. This composite kernel is partially inspired by the kernel function in [6]. The GP-based source codes are partially supported by GPytorch [15]. To optimize the acquisition function, we employed Adam optimizer [25] and the learning rate is set to 0.01.
C.3 Data Collection for Generative Models
To ensure the conditional generative model learns a high-fidelity representation of the Pareto manifold, we employ an adaptive data collection strategy. This method extracts high-quality solutions from the complete evaluation history while dynamically adjusting the resolution of preference vectors based on the complexity of the current approximate Pareto front.
First, we consolidate the cumulative dataset of evaluated solutions from the optimization history to form a candidate pool, denoted as $\mathcal{D}_{pool}$. To determine the appropriate resolution for the scalarization weight vectors, we perform non-dominated sorting on $\mathcal{D}_{pool}$ to identify the non-dominated set (Rank 1, denoted as $\mathcal{F}_{1}$) and the second front (Rank 2, denoted as $\mathcal{F}_{2}$, the non-dominated set with Rank 1 solutions removed). We calculate the combined size of these high-quality fronts as $N_{front}=|\mathcal{F}_{1}|+|\mathcal{F}_{2}|$.
Next, we adaptively determine the number of uniformly sampled scalarization weight vectors, $W$, expected to adequately cover the objective space. The value of $W$ is derived from the front size $N_{front}$ and the objective dimension $M$, following the heuristic $W=(N_{front})^{M-1}$. Therefore, we increase the number of scalarization directions for a higher-dimensional objective space. We then uniformly sample $W$ weight vectors $\{\lambda_{i}\}_{i=1}^{W}$ from the simplex in the objective space.
Finally, we construct the training dataset $\mathcal{D}^{(gen)}$ by scalarizing the entire candidate pool $\mathcal{D}_{pool}$. For each sampled weight vector $\lambda_{i}$, we compute the scalarized values of all solutions in $\mathcal{D}_{pool}$ and extract the top $Q\%$ elite solutions. This strategy is expected to make the generative model trained on a distribution that is performance-oriented via elite selection based on scalarized utility and can also diversely spread across the manifold through the adaptive resolution of $W$ directions.
C.4 Other Detailed Settings
In this work, to convert multi-objective vectors into single-objective scalar values, we instantiate $\mathfrak{s}_{\boldsymbol{\lambda}}(\cdot)$ for all methods using a recently proposed hypervolume scalarization function [59]. In the alternating optimization loop, the number of the sampled candidate solutions via the generative model per task $N_{gen}$ is set to 10,000, which is a common setting as parallel works blending inverse model and expensive multi-objective optimization [52].
Appendix D Detailed Experimental Results
D.1 Detailed Average Hypervolume Comparison
Due to space constraints in the main text, we provide the detailed optimization results in Table 5. The results demonstrate that our proposed generative frameworks, PMT-MOBO-VAE and PMT-MOBO-DDPM, generally achieve superior convergence and solution quality compared to state-of-the-art baselines across distinct tasks.
We first compare our method against single-task approaches, including standard MOBO methods (ST-MOBO, qPAREGO, qEHVI) and Pareto Set Learning variants (PSL-MOBO). As indicated by the statistical tests, our methods significantly outperform standard baselines like ST-MOBO and qPAREGO across all scenarios (marked with ). While the advanced acquisition-based method qEHVI remains competitive on specific landscapes, achieving the top performance on the MAGNETIC problem, PMT-MOBO-DDPM surpasses it on 6 out of 7 benchmarks. This suggests that while specialized acquisition functions are powerful, the generative modeling of the solution manifold offers more consistent performance across diverse problem structures.
We further evaluate the benefit of our generative strategy against existing multi-task optimizers, specifically OCMOBO and PPSL-MOBO, as well as our own ablation PMT-MOBO (without generative sampling). PMT-MOBO-DDPM consistently outperforms OCMOBO and the pure acquisition ablation (PMT-MOBO), confirming that simply utilizing data pooling is insufficient for complex manifolds. The addition of the generative proposal distribution proves critical for navigating high-dimensional or multimodal spaces. Against PPSL-MOBO, a strong multi-task baseline that performed well on DTLZ-3, our DDPM variant achieves higher hypervolume on the majority of real-world tasks (LAMP, SOLAR, UAV), highlighting the superior flexibility of VAE models and diffusion models in capturing complex Pareto geometries compared to polynomial expansions.
A critical comparison is against CDM-PSL, which also utilizes a diffusion model but operates in a single-task setting. The results show a substantial performance gap, with PMT-MOBO-DDPM significantly outperforming CDM-PSL on every benchmark (e.g., achieving $3.37\text{e}4$ versus $1.59\text{e}4$ on DTLZ-1). This validates our core hypothesis regarding inter-task synergy: by pooling data across related tasks to train a conditional generative model, we overcome the data scarcity that cripples single-task deep generative models like CDM-PSL, allowing our optimizer to learn robust inverse maps even under tight evaluation budgets. In summary, PMT-MOBO-DDPM demonstrates the most robust generalization capabilities, achieving the best average Hypervolume on 6 out of 7 benchmarks.
D.2 Ablation Studies
Tables 3 and 4 present the ablation results for VAE and DDPM integrations, respectively. We isolate the impact of the alternating loop by comparing the full method against two decoupled baselines: pure acquisition search (PMT-MOBO w/o GENERATIVE) and pure generative sampling (PLAIN-VAE/DDPM). The results for the VAE variant (Table 3) clearly illustrate the limitations of isolated components. PLAIN-VAE exhibits a high standard deviation across benchmarks (e.g., $7.56\text{e}3$ on DTLZ-1), indicating that while pure generative sampling can occasionally "get lucky" and hit high-quality solutions, it suffers from inherent stochasticity during the optimization process and lacks the precision to consistently refine the Pareto front. Conversely, pure acquisition search often converges prematurely to local optima, failing to explore promising but unseen regions of the design space. By coupling these components, PMT-MOBO-VAE leverages generative sampling to enhance candidate diversity while utilizing the alternating acquisition search to lower the variance of pure sampling. This synergy results in statistically superior convergence and stability across all tested problems.
D.3 Sensitivity Analysis
@@ D.4 Computational Overhead
D.5 Pareto Fronts Generated via Inverse Models
In this subsection, we showcase those exemplar Pareto fronts generated by the inverse models obtained by the algorithm PSL-MOBO in Figure 4, CDM-PSL in Figure 6, PPSL-MOBO in Figure 5, PMT-MOBO-VAE in Figure 7, and PMT-MOBO-DDPM in Figure 8.
Appendix E Workflow of ST-MOBO and PMT-MOBO
Due to the page limit, we detaile the comparative workflow of ST-MOBO and PMT-MOBO here in Algorithm 2.
Algorithm 2 Workflow of ST-MOBO and PMT-MOBO
Output: Final datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$
1:for $t=1$ to $T$ do
2:  Sample preference vector $\boldsymbol{\lambda}\sim\Lambda$
3:  for each task $\boldsymbol{\theta}_{k}$ do
4:   if $\mathcal{A}=\texttt{ST-MOBO}$ then
5:    Select $\mathbf{x}_{k}^{(t)}$ using (4) and (6)
6:   else
7:    Select $\mathbf{x}_{k}^{(t)}$ using (5) and (7)
8:   end if
10:  end for
11:  for each objective $f_{m}$ do
12:   if $\mathcal{A}=\texttt{ST-MOBO}$ then
13:    for each task $\boldsymbol{\theta}_{k}$ do
15:     Train $\mathcal{GP}_{m}$ on $\mathcal{D}_{k}^{(m)}$
16:    end for
17:   else
19:    Train $\mathcal{GP}_{TA,m}$ on $\mathcal{D}_{\text{joint}}^{(m)}$
20:   end if
21:  end for
22:end for
23:return $\{\mathcal{D}_{k}\}_{k=1}^{K}$
Figure 4: Approximated Pareto Front for unseen EMOP tasks in solar rooftop generation (top) and lamp generation (bottom) by PSL-MOBO.
Figure 5: Approximated Pareto Front for unseen EMOP tasks in solar rooftop generation (top) and lamp generation (bottom) by PPSL-MOBO.
Figure 6: Approximated Pareto Front for unseen EMOP tasks in solar rooftop generation (top) and lamp generation (bottom) by CDM-PSL.
Figure 7: Approximated Pareto Front for unseen EMOP tasks in solar rooftop generation (top) and lamp generation (bottom) by PMTO-MOBO-VAE.
Figure 8: Approximated Pareto Front for unseen EMOP tasks in solar rooftop generation (top) and lamp generation (bottom) by PMT-MOBO-DDPM.